# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v1

@@ II Graph Neural Network
In this section, we address Open Questions 1 and 2. We commence by defining the fundamental elements of graph structures, distinguishing between types such as directed vs. undirected and homogeneous vs. heterogeneous graphs. We then provide an overview of different GNN models, including Graph Convolutional Networks (GCNs) [10], Graph Attention Networks [11] (GATs), Graph Autoencoders (GAEs) [105], and Graph Spatial-Temporal Networks [106] (GSTNs), and discuss how they are designed to tackle specific challenges within graph-structured data. Following this, we dive into Open Question 2, where we explore the advantages of node, edge, and graph-level tasks in optimizing the wireless network performance. We offer practical examples of how these task levels can be applied to problems like power allocation, user association, and network deployment in NG-IoT networks. Finally, we present simulation results that compare the performance of different GNN architectures, including the introduction of a hybrid quantum GNN model, offering insights into its potential for enhancing wireless communication systems.
Open Question 1: How Do GNN Paradigms Represent and Process Graph-structured Data Across Different Types of Graphs?
@@ II-1 Background:
    Figure 3: Cellular network as a graph.
    In the context of Graph Neural Networks (GNNs) applied to wireless networks, understanding the fundamental structure of graphs is crucial. A graph is a mathematical representation consisting of a set of vertices (also called nodes) and edges (also called links) that connect pairs of vertices. In general, a graph is represented by a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The set of $|\mathcal{V}|$ vertices is denoted by $\mathcal{V}=\{1,...,v,...|\mathcal{V}|\}$, while the set of edges is defined based on the graph types as follows:
    • Undirected graph: An undirected graph is represented by the set of edges =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$. Here, the edge $(i,j)$ represents an undirected connection between vertices $i$ and $j$.
    • Directed graph: A directed graph, on the other hand, is represented by =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ is the edge directed from vertex $i$ to vertex $j$.
    The illustrations of directed and undirected graphs are represented in the left sub-figure of Fig. 2. Graphs can also be categorized into two types based on the nature of their nodes and edges:
    • Homogeneous graph: These graphs consist of a single type of node and a single type of edge. This simplicity allows for straightforward analysis and processing.
    • Heterogeneous graph: These graphs include multiple types of nodes and edges, as illustrated in the right sub-figure of Fig. 2. This diversity better represents complex real-world networks, such as wireless communication systems, where different devices and connections exist.
    The adjacency matrix of the graph is represented as $\mathbf{A}\in\{0;1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}=1$ if $e_{i,j}\in\mathcal{E}$. For an undirected graph, $\mathbf{A}$ is symmetric, while for a directed graph, it may not be. The node feature matrix $\mathbf{X}$ is a $|\mathcal{V}|\times F_{n}$ matrix, where each row corresponds to a vector of features for a node. Similarly, the edge feature matrix $\mathbf{E}$ is a ||×$|\mathcal{E}|\times F_{e}$ matrix, where each row corresponds to features of an edge. By storing information in the nodes and edges, the graph can capture the complexity of real-world networks, as illustrated in the left sub-figure of Fig. 3. For example, a cellular network can be represented as a graph, where the communication link between a base station and a user corresponds to a node. Edges are formed between all pairs of nodes to represent the interference between different communication links. The feature of the $i$-th node will be the channel gain between the base station and the $i$-th user, denoted as $v_{i}=[h_{i}]$. Meanwhile, $e_{ij}=[h_{i},h_{j}]$ represents the features of the edge between nodes $i$ and $j$, including the interference channels between these links.
  • Graph Neural Network Paradigms
@@
    GNN is a specialized neural network designed for processing and analyzing graph data. GNNs are effective at extracting deep-level topological information, unveiling critical and intricate data characteristics, and enabling efficient data processing. The core idea of GNNs is to learn a mapping function to generate node, edge, or graph representations, known as embedding vectors, based on initial graph information.
@@
    For example, consider a social network graph where each node represents a user, and each edge represents a connection between users. Initially, each node (user) may only contain basic information like user interests. As the GNN processes the graph, the embedding vector of each node is iteratively updated by combining its own information with information from neighboring nodes, as depicted in the graph convolutional layer block in Fig. 4. This iterative process, called message passing, aggregates information from neighboring nodes and edges to update the representation of each node in the graph. Ultimately, each node and edge obtains an embedding that captures the broader network context. The GNN model then processes these node and edge representations to produce node-level, edge-level, or graph-level features for tasks such as node clustering, link prediction, or graph classification. Fig. 4 illustrates the general pipeline of a GNN model, comprising input data, graph convolutional layers, and a downstream task layer.
@@
    GNNs can be broadly classified into different types based on their architecture and function, including Graph Convolutional Networks (GCNs) [10], Graph Attention Networks (GATs) [11], Graph Autoencoders (GAEs) [105], and Graph Spatial-Temporal Networks (GSTNs) [106]. The GCNs use convolution operations directly applied to the graph, updating node representations by aggregating features from neighboring nodes. This method is efficient and scalable, making it eminently suitable for large-scale graph data. Meanwhile, GATs introduce attention mechanisms into GNNs by dynamically assigning weights to the edges during the aggregation process. This approach allows the network to focus on the most relevant parts of the graph structure and enhances the model’s ability to capture complex relationships and dependencies. The attention mechanism ensures that crucial nodes and edges have a greater influence on the final node representation, allowing the GNN to effectively capture complex dependencies within the graph. GAEs are designed for unsupervised learning on graphs, using an encoder to generate lower-dimensional representations and a decoder to reconstruct the graph. They excel in tasks like graph reconstruction and anomaly detection [107]. GSTNs extend GNNs to process time-evolving data by integrating spatial and temporal information, making them suitable for tasks like traffic prediction [108, 109] and dynamic network analysis [110].
@@
    In addition, the integration of GNNs with Reinforcement Learning (RL) has opened new avenues for solving complex decision-making problems that involve graph-structured data [111, 112]. For instance, in this hybrid approach, GNNs encode the graph structure and extract features, which are then fed into RL algorithms to make sequential decisions. This combination is particularly effective in graph-representable scenarios, for example, in network routing [113], robotics [114, 115], and multi-agent systems [116].
@@ II-3 Challenges and future directions:
Open Question 2: How Do Different Task Levels (Node, Edge, and Graph) Contribute to Solving NG-IoT Network Problems?
Figure 5: Node-, edge-, and graph-level tasks.
@@ II-1 Background:
The NG-IoT network problems can be modeled by three distinct task levels as in Fig. 5, including node-, edge-, and graph-levels. Each level represents different aspects of the network and it is associated with specific types of tasks. For example, as illustrated in Fig. 6, an IoT network having multiple access points and serving several devices can be modeled at the node level to focus on tasks like power allocation for individual devices. Edge-level tasks, on the other hand, concentrate on relationships between pairs of nodes, such as user association or channel assignment. Graph-level tasks consider the entire network, addressing challenges like network slicing. Understanding how to formulate wireless problems at the node, edge, or graph level is key to effective problem-solving.
Figure 6: Representation of different levels of graph perspective: (a) The original IoT network, (b) Node-level representation, (c) Edge-level representation, and (d) Graph-level representation

### v2

@@ II Foundations of GNNs for NG-IoT
In this section, we address Open Questions 1, 2, and 3. We commence by introducing the fundamental concept of graphs, distinguishing between key types such as directed vs. undirected and homogeneous vs. heterogeneous graphs. We then provide an overview of Graph Neural Networks (GNNs), including their working principles and prominent architectures, such as Graph Convolutional Networks (GCNs) [10], Graph Attention Networks (GATs) [11], Graph Autoencoders (GAEs) [126], and Graph Spatial-Temporal Networks (GSTNs) [127]. Each of these models is designed to address specific challenges associated with learning from graph-structured data. Following this, we explore Open Question 2 by examining why GNNs are particularly suitable for NG-IoT networks. We analyze the unique characteristics of NG-IoT environments and illustrate how these systems can be effectively modeled as graphs. We then bridge NG-IoT traits to graph-based modeling through the factor graphs inspired by LDPC decoding. GNNs, as learnable generalizations of belief propagation, enable convergence analysis via density evolution (DE) and offer training insights via the extrinsic information transfer (EXIT) charts, reinforcing their suitability for dynamic, scalable, and structured NG-IoT environments. The advantages of node-, edge-, and graph-level tasks are discussed in the context of optimizing key wireless functions, such as power allocation, user association, and network deployment, with practical examples provided. Subsequently, we present simulation results comparing the performance of different GNN architectures, including a proposed hybrid quantum GNN model, which demonstrates promising potential for enhancing wireless communication systems. Finally, in addressing Open Question 3, we investigate the challenges of deploying GNNs in real-world NG-IoT environments. This includes considerations for model implementation, multi-modal data processing, and socio-economic dimensions of GNN. Therein, we explore various types of costs, including computational, energy, and infrastructure costs, that must be considered to ensure cost-effective and scalable GNN adoption in NG-IoT environments. Therein, we analyze key cost considerations, including deployment, operational, and infrastructure expenses, as well as the sustainability challenges associated with deploying GNNs in NG-IoT environments.
Open Question 1: What Are Graphs and Graph Neural Networks?
@@ II-1 Background:
Figure 3: Examples of graphical data representations from various fields. (a) Communication system. (b) Molecules. (c) Social network.
A graph is a mathematical representation consisting of a set of vertices (also called nodes) and edges (also called links) that connect pairs of vertices. In general, a graph is represented by a tuple $\mathcal{G}=(\mathcal{V},\mathcal{E})$. The set of $|\mathcal{V}|$ vertices is denoted by $\mathcal{V}=\{1,...,v,...|\mathcal{V}|\}$, while the set of edges is defined based on the graph types: Undirected graph: =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ represents an undirected connection between vertices $i$ and $j$; Directed graph: =$\mathcal{E}=\{(i,j)|i,j\in\mathcal{V},i\neq j\}$, where $(i,j)$ is the edge directed from vertex $i$ to vertex $j$. Graphs can also be categorized into two types based on the nature of their nodes and edges: Homogeneous graph: Consists of a single type of node and a single type of edge; Heterogeneous graph: These graphs include multiple types of nodes and edges. The illustrations of different types of graphs are represented in Fig. 2. This diversity better represents complex real-world networks, such as wireless communication systems, where different devices and connections exist. The adjacency matrix of the graph is represented as $\mathbf{A}\in\{0;1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}=1$ if ${(i,j)}\in\mathcal{E}$. For an undirected graph, $\mathbf{A}$ is symmetric, while for a directed graph, it may not be. The node feature matrix $\mathbf{X}$ is a $|\mathcal{V}|\times F_{n}$ matrix, where each row corresponds to a vector of features for a node. Similarly, the edge feature matrix $\mathbf{E}$ is a ||×$|\mathcal{E}|\times F_{e}$ matrix, where each row corresponds to features of an edge. By storing information in the nodes and edges, the graph can capture the complexity of real-world networks, as illustrated in Fig. 3. In Fig. 3(a), a hierarchical wireless communication network is represented as a graph, where each system entity, including user equipment (UEs), unmanned aerial vehicles (UAVs), and access points (APs), is modeled as a node. Communication links between entities are depicted as edges, with different edge types capturing distinct connection types. For example, the edge between an AP and the $i$-th UAV is denoted as $e^{\textrm{U-A}}{1}=[g_{i}]$, while the edge between the $i$-th UAV and the $j$-th UE is denoted as $e^{\textrm{U-U}}{ij}=[h_{ij}]$, representing their respective channel gains. Fig. 3(b) illustrates another graph-based example: a chemical structure. Here, atoms such as Carbon (C), Oxygen (O), Nitrogen (N), and Sulfur (S) are represented as nodes, while chemical bonds (single, double, or aromatic) are represented as undirected edges. Fig. 3(c) shows graph modeling in social networks, where individuals are nodes and their interactions, such as friendships, follows, or message exchanges, form the edges. These edges can be directed or undirected, depending on the nature of the relationship. For instance, a “follow” on platforms like Instagram or Twitter is a directed edge, while a mutual friendship results in an undirected edge. Such graph representations are widely used in applications like recommendation systems, influence propagation, and community detection.
@@
  • What is GNNs?: GNN is a specialized neural network designed for processing and analyzing graph data. GNNs are effective at extracting deep-level topological information, unveiling critical and intricate data characteristics, and enabling efficient data processing [8]. The core idea of GNNs is to learn a mapping function to generate node, edge, or graph representations, known as embedding vectors, based on initial graph information.
@@
  • How GNNs works?: As the GNN processes the graph, each node’s embedding vector is iteratively updated by aggregating its own features along with information from neighboring nodes and edges, as shown in the graph convolutional layer block in Fig. 4. In contrast, MLP treats each node independently without leveraging structural relationships, aggregating information only from fully connected layers, while CNN applies local filters to structured grid-based data, capturing spatial dependencies within a fixed neighborhood but lacking adaptability to irregular graph structures. Ultimately, each node and edge obtains an embedding that captures the broader network context. The GNN model then processes these node and edge representations to produce node-level, edge-level, or graph-level features for tasks such as node clustering, link prediction, or graph classification. Fig. 4 illustrates the general pipeline of a GNN model, comprising input data, graph convolutional layers, and a downstream task layer.
@@
  • Types of GNNs and their functionality: GNNs can be categorized based on architecture and function, including Graph Convolutional Networks (GCNs) [10], Graph Attention Networks (GATs) [11], Graph Autoencoders (GAEs) [126], and Graph Spatial-Temporal Networks (GSTNs) [127]. GCNs aggregate node features via convolution operations, making them scalable for large graphs. GATs introduce attention mechanisms to assign dynamic weights to edges, enhancing the model’s ability to capture complex relationships. GAEs, designed for unsupervised learning, use an encoder-decoder framework for graph reconstruction and anomaly detection [128]. GSTNs integrate spatial and temporal information, making them ideal for time-evolving data like traffic prediction [129, 130] and dynamic network analysis [131].
@@
    GNNs also integrate with Reinforcement Learning (RL) for decision-making in graph-structured problems [132, 133]. In such hybrid models, GNNs extract graph features, which RL algorithms leverage for sequential decision-making. This synergy is particularly effective in applications such as network routing [134], robotics [135, 136], and multi-agent systems [137].
@@ II-3 Challenges and future directions:
Open Question 2: Why GNNs for NG-IoT Networks? How Do We Model NG-IoT Systems as Graphs?
Figure 5: Graph representation of some NG-IoT networks: (a) Centralized cellular network, (b) Decentralized D2D network, (c) MEC-enable network, (d) Integrated satellite and terrestrial network.
@@ II-1 Background:
Why GNNs for NG-IoT networks?:
  • i)
    Real-World Challenges of NG-IoT Networks: NG-IoT environments are marked by high complexity and scale, which traditional optimization and deep learning approaches struggle to handle their key characteristics include:
    • Dynamic topology: Devices frequently move, join, or leave the network, leading to constantly changing connectivity and traffic patterns.
    • Heterogeneous Architecture: NG-IoT integrates diverse components such as IoT sensors, mobile devices, MEC servers, RIS, base stations, UAVs, and satellites into a unified but multi-layered network.
    • Non-Euclidean Data: Communication signals and traffic flow originate from complex, irregular domains that do not conform to grid-like structures (e.g., sensor graphs or channel states).
    • Massive Scale and Connectivity: NG-IoT supports ultra-dense device deployments, resulting in graphs having large numbers of nodes and edges.
    • Interdependent Operations: Tasks such as routing, resource allocation, and scheduling rely heavily on node-to-node dependencies and on the contextual awareness of the entire system.
    To address the complexities of NG-IoT networks, modeling and optimization solutions must satisfy critical requirements. Models must be capable of capturing both local and global dependencies across heterogeneous network elements to reflect the interactions in the system. Solutions must handle the dynamically evolving topologies of NG-IoT networks. Given the massive scale of IoT deployments and limited computational capability at the edge, scalability and efficiency are non-trivial. The solutions should remain tractable while maintaining performance across large, distributed graphs. Moreover, context-aware processing is needed to integrate real-time information, such as traffic load, latency, and device states, into the decision-making process. Finally, the models must exhibit strong generalization to unseen scenarios, as NG-IoT systems regularly face novel configurations with new devices and unpredictable traffic behaviors.
  • ii)
    NG-IoT as a Dynamic Factor Graph: The aforementioned characteristics inherently suggest that NG-IoT systems should be modeled as dynamic graphs, where learning must be distributed, adaptive, and structure-aware. To formalize this, we draw inspiration from factor graphs, a structured inference framework used in error correction coding, such as low-density parity-check (LDPC) codes [144, 145, 146, 147]. In LDPC decoding, a bipartite factor graph is constructed with variable nodes representing individual bits of a codeword and factor nodes enforcing parity-check constraints. The decoding process involves iterative message passing among these nodes. Each variable node updates its belief (e.g., log-likelihood ratio) based on incoming messages from neighboring factor nodes, while each factor node enforces local constraints by aggregating the messages it receives and sending back more reliable feedback. This results in a back-and-forth iterative refinement of estimates that converges under well-defined conditions, hence enabling accurate signal recovery even under hostile impairments. This message-passing and constraint-checking mechanism mirrors the interactions in NG-IoT systems, where we have:
    • $\circ$
      Variable nodes: Network entities (e.g., devices, satellites, APs) represent states (e.g., power, queue length, channel quality);
    • $\circ$
      Factor nodes: Network constraints like maximum transmission power or network-wide interference budgets;
    • $\circ$
      Edges: Logical or physical relationships, for example, data dependencies, interference coupling, or task offloading chains.
    However, in contrast to LDPC codes, the topology here is time-varying, with new devices joining and existing ones disconnecting dynamically. This dynamic graph structure naturally aligns with the types of graphs processed by GNNs. Moreover, edges may represent heterogeneous relationships, such as line-of-sight channels, virtual task links, or shared computational dependencies, reinforcing the need for multi-relational and adaptive modeling.
  • iii)
    GNNs as generalized LDPC decoding: The LDPC codes and modern GNNs share a common foundation, message-passing on graphs, but take fundamentally different approaches. The LDPC’s fixed message passing:
    • $\circ$
      Variable-to-Factor: Each device combines messages from all connected constraints using predefined rules.
    • $\circ$
      Factor-to-Variable: Constraints compute responses using exact, hand-crafted formulas.
    Meanwhile, GNNs’ learned message-passing as
    • $\circ$
      Neural message functions: Devices and constraints learn optimal communication strategies through training.
    • $\circ$
      Dynamically weights inputs based on their importance and network conditions.
    Figure 6: Bridging LDPC decoding and GNN-based learning in NG-IoT networks. (a) LDPC-style factor graph representation of NG-IoT systems with fixed message-passing between device states and network constraints. (b) GNN-based generalization using learnable message passing over a dynamic, heterogeneous graph with implicit constraint handling.
    While the belief propagation operates on static factor graphs using fixed update rules such as the sum-product algorithm, GNNs extend this paradigm by introducing learnable message-passing mechanisms tailored for dynamic and heterogeneous graphs like those found in NG-IoT networks. The update process is similar in spirit, but improved flexibility and adaptiveness. Fig. 6(a) illustrates a traditional LDPC-inspired factor graph, where variable nodes (e.g., devices, satellites, UAVs, APs) interact with check nodes (e.g., QoS, power, and bandwidth constraints) through fixed message-passing rules. Fig. 6(b) presents a GNN-based generalization, where diverse node types exchange learnable messages over a dynamic, multi-relational graph. Constraint enforcement becomes soft as well as adaptive, and message updates are learned rather than predefined. This transition highlights how GNNs extend the structure and harness the convergence principles of LDPC decoding to meet the scalability, heterogeneity, and flexibility demands of next-generation IoT systems. Table V presents a structured comparison between traditional LDPC decoding and GNN-based processing in NG-IoT networks. This duality highlights how GNNs generalize the belief propagation framework used in LDPC codes by replacing fixed update rules with learned message functions, enabling dynamic, data-driven processing over heterogeneous and evolving network topologies. The table outlines key analogies across graph types, node and edge semantics, message-passing mechanisms, convergence tools, and application domains, thus bridging well-established coding theory with modern GNN-based learning for IoT.
    TABLE V: GNN-LDPC Duality in NG-IoT Systems
    Aspect LDPC Decoding GNNs in NG-IoT Networks
    Graph Type Static bipartite factor graph (variables $\leftrightarrow$ parity checks) Dynamic multigraph (devices, APs, servers, constraints, etc.)
    Variable Nodes Bits (e.g., codeword elements $x_{i}$) Device states (e.g., power, location, queue length, CSI)
    check nodes Parity-check constraints Physical/logical constraints (e.g., interference limits, delay, energy budgets)
    Edges Binary connections between variables and checks Typed edges for wireless links, task dependencies, or shared resource constraints
    Message Type Log-likelihood ratios (LLRs) Learnable embeddings
    Message Passing Sum-product algorithm (fixed) Trainable functions
    Learning No learning (hand-crafted rules) Data-driven learning via backpropagation
    Convergence Tool Density evolution (analytical) Neural density evolution [148]; convergence under Lipschitz conditions
    Information Flow EXIT charts for decoding visualization EXIT-inspired metrics to diagnose layer-wise training dynamics
    Applications Error correction (bit decoding) Power control, scheduling, routing, anomaly detection in NG-IoT
  • iv)
    Convergence analysis via density evolution and EXIT charts for learning analysis:
    • $\circ$
      Density evolution (DE) [144]: LDPC’s DE tracks message distributions across iterations to predict decoding thresholds. For GNNs, Matthieu et al. [148] has shown that, under mild assumptions (e.g., Lipschitz continuity), GNN message-passing admits a state-space evolution analogous to DE. This shows that GNN updates can be viewed as stochastic operators acting over a structured graph domain, mirroring DE in LDPC systems. Therefore, the iterative message-passing process converges to a fixed-point embedding. For instance, DE can help evaluate the stability and performance of GNN-based algorithms in resource allocation or routing under dynamic conditions.
    • $\circ$
      EXIT (Extrinsic Information Transfer) charts: EXIT charts [147] constitute another core concept in LDPC decoding, used to visualize how mutual information between variable and check nodes evolves during iterations. In the GNN context, such visualization is not yet standard, but analogous metrics have been explored. For instance, mutual information tracking between hidden representations across layers has been proposed in [149], where Kenta et al. showed how deeper GNNs suffer from representation collapse. By adapting EXIT-style visual tools to plot mutual information layer by layer, practitioners can better diagnose issues like over-smoothing or underfitting, and design more robust and interpretable GNNs for NG-IoT applications. Furthermore, they can also provide tangible design guidelines, concerning the best converging activation order of handling the nodes’ inputs. They can also be used for mitigating the signaling overhead by identifying the nodes in the immediate neighborhood that are capable of expediting convergence, while neglecting those providing marginal improvements.
    • $\circ$
      Bridging insight: Although DE and EXIT charts originated from coding theory, they provide a solid analytical foundation for understanding the GNN behavior. GNNs can be viewed as learnable generalizations of belief propagation, where instead of fixed message-update rules, the aggregation and update functions are parameterized by neural networks. This flexibility introduces challenges in theoretical analysis but also allows tools from LDPC research to be reinterpreted to provide convergence guarantees and learning diagnostics for GNN-based solutions. To elaborate a little further from an LDPC perspective code having short cycles in their factor-graph invariably perform poorly. Hence further research is required for eliminating short cycles in the GNNs. By the same token, LDPC codes having rates between 1/3 and 2/3 tend to have the best performance. Hence the factor-graph-based design of GNNs has a high promise for future research.
Given the aforementioned characteristics and analytical findings, GNNs offer a naturally compatible and theoretically grounded framework for NG-IoT networks:
  • Structure-aware modeling: GNNs operate directly on graph-structured data, allowing them to represent complex relationships among heterogeneous network entities (e.g., devices, APs, UAVs, satellites).
  • Dynamic message passing: Unlike static models, GNNs use learned message-passing mechanisms that adapt to evolving network topologies and conditions, enabling real-time, context-aware, and distributed decision-making.
  • Theoretical grounding: By generalizing belief propagation from LDPC codes, GNNs can leverage insights from density evolution and EXIT charts to analyze convergence behavior and guide the design of stable, efficient architectures.
  • Scalability and generalization: GNNs naturally generalize across varying graph sizes and structures, making them robust to unseen network conditions. They remain scalable through techniques such as graph sampling, partitioning, and hierarchical modeling.
  • Training diagnostics and interpretability: EXIT charts provides mutual information tracking that helps identify and mitigate issues like over-smoothing and representation collapse in deep GNNs, improving learning efficiency and interpretability.
Graph representation of NG-IoT networks:
To address NG-IoT network challenges effectively, they must first be formulated as graph-based problems. To do so, this involves key steps:
  1. 1.
    Identifying graph components: Define key elements for the NG-IoT network and map them to graph components.
    • Nodes: User equipment, base stations, access points, UAVs, satellites, MEC servers, RIS, etc.
    • Edges: Communication links, interference channel, data dependencies.
    • Graph types: Directed vs. undirected, homogenous vs. heterogeneous graphs.
  2. 2.
    Mapping network data to graph information: Associating network data with the corresponding graph elements.
    • Node features: Device type, mobility, energy consumption, transmission power, individual data.
    • Edge features: Link quality, latency, interference level, channel gain.
    • Global network characteristics such as connectivity density or overall traffic load.
  3. 3.
    Graph construction based on network characteristics
    • Centralized Networks (e.g., Cellular Networks)
      • Description: Rely on centralized infrastructure like base stations for communication.
      • Key Features: Wide coverage, high reliability, scalable.
      • Graph Representation:
        • i)
          Vertices: Base stations, IoT devices, and core network elements.
        • ii)
          Edges: Links between devices and base stations and base stations to the core.
      • Example: Fig. 5(a) depicts a typical cellular system where UEs connect to APs, and APs forward data to a central server. This forms a heterogeneous graph with nodes representing UEs, APs, and the server, and edges modeling wireless and wired communication links. Such structure supports centralized resource control and coordination.
    • Decentralized Networks (e.g., D2D, Mesh/Ad-hoc Networks)
      • Description: Enable direct device communication without centralized infrastructure.
      • Key Features: Low latency, energy-efficient, resilient.
      • Graph Representation:
        • i)
          Vertices: IoT devices.
        • ii)
          Edges: Direct, dynamic links between devices.
      • Example: Fig. 5(b) shows devices communicating directly without central infrastructure. Each device is a node, and edges represent dynamic, peer-to-peer links. This homogeneous graph structure captures local interactions in energy-efficient and latency-sensitive applications.
    • Edge-Enabled Networks (e.g., MEC-enabled Networks)
      • Description: Use edge computing for localized data processing.
      • Key Features: Low latency, real-time analytics, AI/ML integration.
      • Graph Representation:
        • i)
          Vertices: IoT devices, edge servers, cloud data centers.
        • ii)
          Edges: Links between devices and edge servers, and edge servers to the cloud.
      • Example: Fig. 5(c) showcases a network architecture enhanced with mobile edge computing, where edge servers serve as intermediaries between end devices and the cloud. The graph includes nodes for IoT devices, edge servers, and cloud infrastructure. Edges connect devices to their respective edge servers and edge servers to the centralized cloud.
    • Hybrid Networks (e.g., Integrated Terrestrial-Satellite Networks)
      • Description: Combine terrestrial and satellite communication for global coverage.
      • Key Features: Ubiquitous connectivity, high reliability.
      • Graph Representation:
        • i)
          Vertices: IoT devices, base stations, satellites, ground stations.
        • ii)
          Edges: Links between terrestrial and satellite components.
      • Example: Fig. 5(d) illustrates an integrated terrestrial infrastructure (e.g., base stations and IoT devices) with aerial and satellite elements (e.g., UAVs, GEO satellites). The graph representation includes heterogeneous nodes, such as, IoT terminals, terrestrial base stations, UAVs, and satellites, with edges capturing both ground and space communication links.
  4. 4.
    Problem formulation using graph structures (Incorporating code, edge, and graph-level tasks)